AnimeGhoul8863:

Let f(x) = x − 2 and g(x) = x2 − 7x − 9. Find f(g(−1)).

AnimeGhoul8863:

@nuts

nuts:

ok, i want you to guess where to start

AnimeGhoul8863:

im alittle confused cause i dont know what to put for x

nuts:

take a wild guess where to start :)

AnimeGhoul8863:

x2 − 7x − 9. (-1)2-7(-1)-9

nuts:

that's the right place to start c: what do you get?

AnimeGhoul8863:

\[-1^2 = -2\] right

nuts:

noo (-1)^2=(-1)(-1)=1

AnimeGhoul8863:

ok so 1^2-7(1)-9

nuts:

no :( \( x^2 − 7x − 9; x=-1\) \((-1)^2-7(-1)-9=?\)

AnimeGhoul8863:

i know but u said -1^2=1so i put \[1-7(1)-9=\]

nuts:

ok slow down there. (-1)^2=1. -1^2=-(1^2)=-1 but, x=-1, so you need to be plugging in -1's for g(x) when you see x

AnimeGhoul8863:

=-1?

nuts:

@ThisGirlPretty

AnimeGhoul8863:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @nuts no :( \( x^2 − 7x − 9; x=-1\) \((-1)^2-7(-1)-9=?\) \(\color{#0cbb34}{\text{End of Quote}}\) (-1)^2-7(-1)-9=-1

AnimeGhoul8863:

am i wrong

AnimeGhoul8863:

brb

AnimeGhoul8863:

Tgp am i wrong